# AI-enabled tuberculosis screening in a high-burden setting using cough sound analysis and speech foundation models

### v1

@@ Discussion
Our study highlights the potential of AI-enabled cough sound analysis as a scalable and effective TB screening tool, in both PLHIV and HIV negative individuals. The model demonstrated strong performance, with improved accuracy when incorporating clinical features, and its robustness when using both laptop and mobile phone recordings suggests promise for point-of-care applications. AI-enabled cough analysis could enhance TB detection, especially in high-risk populations and resource-limited settings. Future research should assess generalisability across regions, extend to subclinical TB, and benchmark performance against chest X-rays or molecular diagnostics. Exploring model compression and training efficiency will also be important for real-world deployment in resource-constrained settings.
Contributors
The study was conceived by MK, MM, NM, and GJB. Data collection and management were carried out by SC, SM, NS, MM, BM, MK, GJB, and NM. Model development was led by NM and BH. Data analysis and interpretation were performed by NM, BH, GJB and MK. NM, MK, and BH wrote the first draft of the manuscript. Literature search was conducted by MK, GJB, and NM. Figures were produced by NM and BM. All authors had full access to all the data in the study and contributed to the writing, review, and approval of the final version of the manuscript.
Declaration of interests
We declare no competing interests.
@@
The TB cough sound dataset and code used in this study will be made available by the University of Sheffield to non-profit researchers upon reasonable request after publication, subject to a data use agreement. Requests should be directed to the corresponding author.
Acknowledgments
The authors would like to thank Ms Regina Banda and Ms Winfrida Mashili for their support with data collection, and Ms Sheba Nalwaba for assistance with data entry. We are also grateful to the management and staff of Kanyama and Chawama First-Level Hospitals for providing a supportive environment for data collection. This work was funded by the UK Higher Education Innovation Fund (X/179090) and the UK Engineering and Physical Sciences Research Council (EPSRC) Impact Acceleration Account (R/185787-11-1).

### v2

@@ 4 Discussion
5 Conclusion
Our study highlights the potential of AI-enabled cough sound analysis as a scalable and effective TB screening tool, in both PLHIV and HIV negative individuals. The model demonstrated strong performance, with improved accuracy when incorporating clinical features, and its robustness when using both laptop and mobile phone recordings suggests promise for point-of-care applications. AI-enabled cough analysis could enhance TB detection, especially in high-risk populations and resource-limited settings.
Future research should assess generalisability across regions, extend to subclinical TB, and benchmark performance against chest X-rays or molecular diagnostics. Exploring model compression and training efficiency will also be important for real-world deployment in resource-constrained settings.
@@ 5 Conclusion
The TB cough sound dataset and associated code used in this study will be made available to qualified academic and non-profit researchers upon reasonable request following publication, subject to a data use agreement with the University of Sheffield. Data access requests should be submitted to the corresponding author.
Acknowledgements
This work was supported by the UK Higher Education Innovation Fund (X/179090) and the UK Engineering and Physical Sciences Research Council (EPSRC) Impact Acceleration Account (R/185787). The authors would like to thank Ms Regina Banda and Ms Winfrida Mashili for their support with data collection, and Ms Sheba Nalwaba for assistance with data entry. We are also grateful to the management and staff of Kanyama and Chawama First-Level Hospitals for providing a supportive environment for data collection.